OpenStudy (anonymous):

there are 1176 students in central high school.there are 50 more girls than boys, how many boys are there

OpenStudy (kamille):

let's say that boys are x, then there are x+50 girls (because there are 50 more girls then boys),so x+x+50=1176 can you solve it for me?

OpenStudy (anonymous):

-588?

OpenStudy (kamille):

how it can be a minus?

OpenStudy (kamille):

what is x+x?

OpenStudy (anonymous):

boys be x and girls be x+50 x+x+50=1176 2x=1176-50 2x=1126 x=1126\2 x=563 boys=563 and girls=563+50=613

OpenStudy (anonymous):

since there are 50 more girls subtract that from the 1176 then take the sum and divide by two

OpenStudy (anonymous):

and the 50 to the sum for the girls